### v2

# Quantum Simulation of Spin-Boson Models with Structured Bath

Abstract
The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems. Trapped ions present a natural platform for simulating the quantum dynamics of such models, thanks to the presence of both high quality internal qubit states and the motional modes of the ions that can simulate the relevant quantum degrees of freedom. In our work, we extend the previous body of work that focused on coherent coupling of the spins and bosons to perform quantum simulations with structured dissipative baths using the motional states of trapped ions. We demonstrate the capability for adjusting the bath’s temperature and continuous spectral density by adding randomness to fully programmable control parameters. Subsequently, we simulate the dynamics of various spin-boson models with noise spectral densities constructed from coupling to several dissipative harmonic oscillator modes. The experimental outcomes closely align with theoretical predictions, indicating successful simulation of open quantum systems using a trapped-ion system.
Interaction of a quantum system with its environment in an open dissipative system can significantly influence the dynamics of the system. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987); Schindler et al. (2022), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al. (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an interesting challenge for quantum simulators Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022).
Trapped ions provide a highly controllable system of spins and motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a handful of dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Jia et al. (2022); Chen et al. (2023a), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Gorman et al. (2018); MacDonell et al. (2023); Whitlow et al. (2023); Valahu et al. (2023). These previous studies did not consider the effects of dissipation, which is often described by the coupling of a system to a continuum of bath modes, that can determine the system’s long-term behavior such as reaching a steady state.
In this work, we perform quantum simulations of spin-boson models with continuous spectral densities using trapped ions and their motional modes. Specifically, we simulate spectral densities composed of several Lorentzian peaks, which is characteristic of baths in light-harvesting protein complexes Olbrich et al. (2011); Chandrasekaran et al. (2015); Lee and Coker (2016); Kim et al. (2018); Tamascelli et al. (2019). Furthermore, we simulate the sub-Ohmic, Ohmic, and super-Ohmic spin-boson models Leggett et al. (1987), which are ubiquitous in various open quantum systems ranging from nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011) and superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023) to molecules in which proton-transfer reactions occur Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). Finally, we simulate the vibrationally-assisted energy transfer (VAET) Gorman et al. (2018) in a simple molecular model where dephasing is added to the vibrational mode. The various spin-state population dynamics, which display behaviors such as damped oscillations, coherent revivals, and relaxation to a steady state, are accurately captured by our quantum-simulation results. We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
Capturing the dissipative behavior of the spin-boson model, as well as tuning the bath’s temperature, is enabled by adding randomness to the control parameters, such as the laser frequency and phase. Remarkably, the impact of these control signals is directly analogous to motional dephasing, the leading source of noise in our experimental setup Kang et al. (2023a). This serves as a lower limit for the level of noise that can be introduced in our quantum simulation process. While the quantum simulations presented here are limited to the weak spin-bath coupling regime, our demonstration of adding or using noise Guimarães et al. (2023); Leppäkangas et al. (2023) in analog quantum simulations is an important ingredient for a potential quantum advantage in simulating open quantum system models that are intractable using classical simulation methods MacDonell et al. (2021); Kang et al. (2024).
Results
Models
The spin-boson model describes a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by
[…]
A spin-boson model with a structured bath is often characterized by a spectral, density $J(\omega)$, composed of multiple Lorentzian peaks, each of which can be represented by a noisy harmonic oscillator Lemmer et al. (2018). Specifically, we consider a spin coupled to dephased oscillators (hereafter referred to as the dephased spin-oscillator model). The single-oscillator case is described by the Hamiltonian
$\hat{H}_{SO}=\hat{H}_{S}+\frac{\kappa}{2}\hat{\sigma}_{Z}\otimes(\hat{b}+\hat{b}^{\dagger})+\nu\hat{b}^{\dagger}\hat{b}$ (2)
and a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}^{\dagger}\hat{b}$, where $\hat{b}$ is the annihilation operator of the oscillator, such that the composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{SO},\rho]+\sum_{j}(\hat{L}_{j}\rho\hat{L}_{j}^{\dagger}-\frac{1}{2}\{\hat{L}_{j}^{\dagger}\hat{L}_{j},\rho\})$ Lindblad (1976). In a reasonable regime of parameters ($\Gamma<\nu/2$, $\beta\ll 2\pi(\Gamma/2)^{-1}$) Lemmer et al. (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
$J_{\rm Lo}(\omega)=\kappa^{2}\Big{(}\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega-\nu)^{2}}-\frac{\Gamma/2}{(\Gamma/2)^{2}+(\omega+\nu)^{2}}\Big{)},$ (3)
where we assume that the oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu)$.
The spectral density of the bath determines the dynamics of the spin to leading order in $\kappa T$, where $T$ is the evolution time Breuer and Petruccione (2002); sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\kappa$ regime, which extends straightforwardly to the model with multiple oscillators. Of particular interest is the unbiased ($\epsilon=0$) spin-boson model Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), where the equilibrium donor-state population is $1/2$ for both dephased spin-oscillator models and spin-boson models. Thus, for weak spin-bath coupling, the population dynamics of the two models show an excellent match at all times; the match in short (long) time scale is achieved due to matching spectral densities (equilibrium populations). Unless specified otherwise, we use $\epsilon=0$; results for $\epsilon\neq 0$ can be found in Fig. 5.
A more commonly considered model is a spin coupled to damped oscillators (hereafter referred to as the damped spin-oscillator model); for example, the single-oscillator case is described by the Hamiltonian $\hat{H}_{SO}$ and a pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma(\bar{n}(\nu)+1)}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma\bar{n}(\nu)}\hat{b}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021). For this model, a stronger result holds: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\kappa T$ Tamascelli et al. (2018); sup . Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al. (2018); Schlawin et al. (2021). In this work, we focus on dephased oscillators to represent the structured bath, which is readily performed using programmable randomized control operations on a single species of trapped ions.
Single oscillator model
The dephased spin-oscillator model can be simulated using trapped ions by applying spin-dependent kick (SDK) operations, which induce the coupling between the trapped-ion qubit and the motional mode (see Methods for details). By adding randomness to the control parameters of the SDK operations, we can also prepare the thermal state and simulate the dephasing of the motional mode, by averaging the results over many random trials. The key idea is that noisy (heated or dephased) evolution, described by the Lindblad master equation, is equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Cai and Barthel (2013); Chenu et al. (2017).
Figure 1: Demonstration of the tunability of the bath’s temperature and spectral linewidth. (a) Dynamics of the coherent spin-oscillator model $\hat{H}_{SO}$ with various values of average phonon number $\bar{n}$ of the oscillator’s initial state. (b) Dynamics of the spin-boson model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ and various values of $\Gamma$. For both (a) and (b), $\epsilon=0$, $\nu=\Delta$, and $\kappa=0.1\Delta$. Lines (dots) represent theoretical predictions (experimental data) and error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times. (c) Revival amplitude versus $\bar{n}$. Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles). (d) Fitted values of $\bar{n}$, where the experimental data are fitted to the predictions of the noise-aware model. Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling. (e) Fitted values of $\Gamma$, where the experimental data are fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$. Solid line depicts $y=x$. In (d) and (e), the blue (brown) shaded region represents the expected uncertainty of the respective value when the populations are averaged over 20 (200) random trials. (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$. The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
We first demonstrate the programmability of the oscillator bath’s temperature, quantified by the average phonon number $\bar{n}$ of the bath oscillator’s initial state. Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion in the motional ground state, which prepares the thermal state when averaged over many random trials (see Methods for details). Figure 1a illustrates the simulated coherent time evolution with respect to $\hat{H}_{SO}$, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with an envelope of slow beat-note that signifies resonant ($\nu=\Delta$) energy exchange between the system and the bath oscillator Eberly et al. (1980). Our initial thermal state is prepared as an ensemble of randomly displaced coherent states. A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$. We see a notable agreement between theoretical predictions (solid lines) and experimental results (dots).
We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup. We use spin-boson models with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma$ [see Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from separate experiments to characterize imperfect cooling and finite motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 1c) and fitted $\bar{n}$ (Fig. 1d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can be used as part of the dissipation model, and thus is not always a bottleneck for the accuracy of our simulations.
Next, we show the tunability of the dephasing strength, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution (see Methods for details). This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020).
Figure 1b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a phase transition from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$). As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime. We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case (see, for example, the phase difference between the underdamped and overdamped case at $t\approx 2\pi/\Delta\times 10$). These features match very well with the theoretical predictions of spin-boson models.
The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters. The shaded areas in Fig. 1d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by averaging the results over either 20 (pink) or 200 (blue) random trials. Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population. The experimental data is consistent with the standard deviation for 20 trials. We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al. (2023); sup .
Engineering bath spectral densities
We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain. The spin-boson model with $J(\omega)$ in Fig. 2b is simulated using the SDK operations over three motional modes. Fig. 2a shows the frequency of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency. The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from the motional mode determines the frequency at which the bath spectral density is considered in Figure 2b. Here, we set the electronic coupling strength $\Delta=500$ cm-1 sup ; Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
Figure 2c shows the results for simulating spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm White and Feiguin (2004); Prior et al. (2010) in the interaction picture Nuomin et al. (2022). The donor population features underdamped coherent oscillations, with contributions from all three modes. The measured population closely follows the theoretical predictions.
The coherence time of the motional modes in our setup of $\sim 5.18(67)$ ms is the leading source of error Wang et al. (2020); Kang et al. (2023a), and is comparable to the experimental time for simulating the time evolution of 2-mode (1.63 ms) and 3-mode (2.56 ms) cases. For these simulations, the contribution of the motional decoherence in the experiment is identical to the random kicks we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al. (2023a); Milne et al. (2021); Keller et al. (2021).
Figure 2: Simulations of molecular energy transfer in model structured baths. (a) Dashed lines represent the radial motional modes strongly coupled by the Raman beam. Black (grey) lines denote the modes with non-zero (zero) coupling strength to the target ion. The three solid lines (shaded regions) depict the mean value (standard deviation) of the randomized pulse frequencies. (b) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$. The dashed curves depict the Lorentzian peak generated by the pulses of corresponding colors in (a). The orange (brown) solid curves illustrate the spectral density summed over two (three) peaks. (c) Time evolution of the spin-boson models with $J(\omega)$ in (b). The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling. Curves (dots) represent theoretical predictions (experimental data). Error bars denote the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
Figure 3: Simulations based on Leggett’s spin-boson model. (a) Spectral densities of the Leggett’s model. Dot-dashed curves are the sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities described by Eq. (4) ($A=0.1$, $\omega_{c}=10\Delta$). Solid curves are the spectral densities simulated in our experiments using two motional modes. These spectral densities are designed to match the dot-dashed curves near $\omega=\Delta$. Insets compare the solid and dot-dashed curves within the interval $\omega\in[0.85\Delta,1.15\Delta]$. (b) (c) (d) Dynamics of the spin-boson model for $s=0.5$, $1.0$, and $2.0$, respectively. The solid curves represent the theoretical predictions of the spin-boson models with spectral densities given by the solid curves in (a). The red dashed curves are expected results, derived from numerical simulations of the density matrix of the qubit and motional modes subject to control operations and hardware noise. Circles denote experimental data and error bars indicate the standard deviation over 60, 20, and 20 trials for (b), (c), and (d), respectively. Each trial is performed with randomly drawn set of parameters and repeated 100 times. (e) Time evolution of the spin-boson models. The solid (dot-dashed) curves show the evolution with spectral densities given by the solid (dot-dashed) curves in (a).
Simulating the Leggett spin-boson models
We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density of the bath is given by
$J_{\rm Legg}(\omega)=A\omega^{s}\omega_{c}^{1-s}e^{-\omega/\omega_{c}},$ (4)
where $\omega_{c}$ is the cutoff frequency, and $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively. This model is widely used to characterize the noise in nanomechanical devices Seoanez et al. (2007); Galland et al. (2008); Chen et al. (2011), superconducting circuits Magazzù et al. (2018); Thorwart et al. (2004); Kaur et al. (2021); Gulácsi and Burkard (2023), and proton-transfer reactions Cukier and Morillo (1989); Ohta et al. (2006); Craig et al. (2007). The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al. (1987); Bulla et al. (2003); Vojta et al. (2005) and classical simulations Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Kast and Ankerhold (2013); Nalbach and Thorwart (2013); Otterpohl et al. (2022); Nacke et al. (2023).
We use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ Meier and Tannor (1999). To find the optimal set of parameters $\kappa$’s, $\Gamma$’s, and $\nu$’s in Eq. (3) that best represent $J_{\rm Legg}(\omega)$, well-established global optimization algorithms Wolz (2022) are used to minimize the cost function, defined as the integral of the absolute difference between $J_{\rm Legg}(\omega)$ and the sum of corresponding Lorentzian spectral densities over a given bandwidth sup .
As a first step, we consider weak spin-bath coupling ($A=0.1$) and designed the bath spectrum for a limited bandwidth near resonance $\omega\in[0.9\Delta,1.1\Delta]$ using 2 Lorentzian peaks. Figure 3a shows the obtained spectra from the modes (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively. The results for the donor population dynamics interacting with a bath featuring these approximated spectral densities are shown in Fig. 3b-d. We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale depending on the level of noise in the bath. The experimental data match well with theoretical predictions, where the deviations in the $s=2$ case result from insufficient number of Trotterization steps sup .
Figure 4: Simulating sub-Ohmic ($s=0.5$) bath with 10 oscillators. (a) Comparison between the sub-Ohmic spectral density (dot-dashed blue) and its approximation constructed by 10 noisy oscillators (solid red). Each dashed line represents the Lorentzian spectral density generated by each of these 10 oscillators. (b) Comparison of the population dynamics generated by the sub-Ohmic spectral density (dot-dashed blue) and its 10-mode approximation (solid red).
Figure 3e compares the predicted population dynamics of the spin-boson model with ideal spectral density $J_{\rm Legg}(\omega)$ and the corresponding approximated spectral density. While the dynamics match very well for the Ohmic and super-Ohmic cases, the sub-Ohmic case shows a non-negligible deviation due to the gap in the two spectral densities within the bandwidth (insets in Fig. 3a). In future experiments with a larger number of motional modes Porras et al. (2008), we expect to match a wider range of spectral densities and thus more accurately simulate the sub-Ohmic spin-boson model, as suggested by the 10-mode numerical example in Fig. 4.
Vibration-assisted energy transfer
Finally, we simulate the VAET model Gorman et al. (2018) with nonzero energy detuning $\epsilon$ between the spin states. Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$. In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition.
Figure 5: Simulations of vibrationally-assisted energy transfer (VAET). We set $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$. (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$. Solid (dashed) curves depict theoretical predictions of the target (noise-aware) model. Circles represent experimental data and error bars are standard deviation over 100 repetitions. In (a), bath temperature is fixed at 0 K ($\bar{n}=0$) and $\nu$ is varied. In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated. (c) Time evolution of the donor-state population for noisy VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K. Solid (dashed) curves depict theoretical predictions of the spin-boson (dephased spin-oscillator) model. Circles denote experimental data. For $\kappa=0\>\text{cm}^{-1}$, error bars indicate the standard deviation over 100 repetitions. For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
Figure 5a shows the experimental data for simulating the coherent spin-oscillator model $\hat{H}_{SO}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$). When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al. (2018). Otherwise, energy remains localized at the donor site. Figure 5b, shows the impact of initial temperature of the oscillators on the resonant VAET ($\nu=104\>\text{cm}^{-1}$). We simulate baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively (see Methods). The energy transfer is faster for higher temperature bath in the beginning, as more vibrational quanta assist the energy transfer. However, high temperature leads to a rapid decay of oscillations over time as the coherence is lost.
In Fig. 5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling. We note that as $\kappa$ is increased, the results deviate more from the theoretical predictions of the damped spin-oscillator model, which is non-perturbatively equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al. (2018); Tamascelli et al. (2018). This is consistent with the expectation that the method of using the dephased spin-oscillator model to approximate the spin-boson model is valid only in the weak-$\kappa$ regime. While the short-time dynamics between the two models match as the 2-point correlation functions are the same, the long-time dynamics deviate due to mismatch in the higher-order correlation functions for higher coupling strengths sup .
Discussions and outlook
Simulations of large-scale spin-boson models with strong coupling in the presence of dissipation is crucial yet challenging problem in physical Caldeira and Leggett (1983); Weiss et al. (1987); Schindler et al. (2022), chemical Garg et al. (1985); Gilmore and McKenzie (2005), and biological Huelga and Plenio (2013); Jang and Mennucci (2018) systems. While classical simulation approaches continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al. (2021); Kang et al. (2024). Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method, a state-of-the-art method for simulating open quantum systems Somoza et al. (2019). For simulating a one-dimensional chain of $N_{s}$ electronic states each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix. A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics. For these types of problems, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales only as $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al. (2024). The current state-of-the-art trapped ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al. (2022); Chen et al. (2023b). Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems can open up a new computational framework in science.
The methods developed in this work utilizes fully programmable control methods to simulate dissipative processes. This approach can readily be applied to existing simulations of chemical processes to include coupling to environment Whitlow et al. (2023); Valahu et al. (2023); Sun et al. (2023); Zhang et al. (2024); Li et al. (2021). The dephased oscillator method can be used to realize non-Gaussian bath models Linowski et al. (2022); Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Duan et al. (2017); Otterpohl et al. (2022) and role of bosonic dephasing in molecular energy transfer Grad et al. (1988); Goldsmith et al. (2006); Plenio and Huelga (2008); Hayes et al. (2010); Moreira et al. (2020); Yoneda et al. (2021). High accuracy simulations will require development of efficient characterization methods Kang et al. (2023b) and novel control protocols in the trapped ion system sup ; Lobser et al. (2023) to implement various interaction terms in the target Hamiltonian. Combined with damped spin-boson models that can be implemented using sympathetic cooling Lemmer et al. (2018); Schlawin et al. (2021), the long-chain trapped ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
During the preparation of our manuscript we became aware of a complementary work So et al. (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode.
Acknowledgements.
The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work. This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No. OMA-2120757 (experimental concepts and validation). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation).
Methods
Experimental implementation
Figure 6: Experimental setup and instruction of operations. (a) Schematic of the experimental setup illustrates the orientation of relevant fields. The magnetic field and the direction of the individual beam is perpendicular to the trap surface, while the global beam propagates is parallel to it. The differential wave vector $\Delta\textbf{k}$ is therefore oriented at a 45-degree angle to the trap surface, aligning with one of the radial axes. The global beam comprises two tones driving the blue (BSB) and red (RSB) sideband transitions, respectively. (b) Simplified energy levels of the 171Yb+ ion and the laser pulses used for implementing the on-resonance spin-dependent kick (SDK) operation. (c) Visualization of the motional-state evolution in phase space when a SDK operation of duration $\tau$ is applied, where the qubit is in the eigenstate of the operator $\hat{\sigma}_{\phi_{s}}\equiv\hat{\sigma}_{+}e^{i\phi_{s}}+\hat{\sigma}_{-}e^{-i\phi_{s}}$. Here, $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit, $\phi_{s}$ ($\phi_{m}$) is the spin (motion) phase of the laser beams, $\tilde{\Omega}$ is the Rabi frequency of the sideband transition, and $\delta_{m}$ is the motion detuning. (d) Instruction of operations for simulating the dephased spin-oscillator model, which approximates the spin-boson model. ‘H’ represents the Hadamard gate on the qubit. ‘Kick (heat)’ represents the SDK operations with random $\phi_{m}$ that prepare the thermal state on average. ‘SQ’ (‘Kick (coh/dep)’) represents the single-qubit rotations (SDK operations) that simulate the first (second) term in Eq. (5). Coherent (dephased) spin-oscillator model is simulated by using a fixed (randomly drawn) $\delta_{m}$ value at each step.
The simulations of spin-boson models are performed using a linear chain of 171Yb+ ions confined in a micro-fabricated surface trap Revelle (2020), as illustrated in Fig. 6a. The spin states are encoded in the hyperfine energy levels in the 2S1/2 manifold (qubit states) as |$\ket{0}\equiv\ket{F=1;m_{F}=0}$ and |$\ket{1}\equiv\ket{F=0;m_{F}=0}$. The bosonic bath modes are encoded as the collective radial motional modes of the ion chain. Details of our experimental setup can be found in Ref. Wang et al. (2020).
Through the sequential application of Doppler cooling, EIT cooling, and sideband cooling techniques, the zigzag motional mode of the ion chain can be cooled to $\bar{n}=0.036(16)$. Following this, a $370$ nm laser is utilized to pump the ions to their ground state |$\ket{1}$. Transitions between the qubit states are performed using stimulated Raman transitions, driven by a pair of laser beams Wang et al. (2020). Acousto-optic modulators are employed to manipulate the frequency and phase of each laser beam, thereby enabling single-qubit gates: we apply a bit flip ($X$) gate to initialize the ion to |$\ket{0}$ state. The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous blue- and red-sideband transitions as shown in Fig. 6b. Figure 6c visualizes how the motional state evolves in phase space when a SDK operation of duration $\tau$, spin (motion) phase $\phi_{s}$ ($\phi_{m}$), motion detuning $\delta_{m}$, and sideband Rabi frequency $\tilde{\Omega}$ is applied Lee et al. (2005). Subsequent to the final operation, the qubit-state population is accessed via the state-dependent detection technique.
We strategically use the center ion of a 7-ion chain, as it exhibits negligible coupling to the 2nd, 4th, and 6th modes (See Fig. 7). This arrangement maximizes the effective spacing between the motional-mode frequencies and therefore suppresses cross-mode coupling (driving modes that are not targeted by the SDK operation).
Figure 7: Ion-Mode coupling strengths of a 7-ion chain. Motional-mode frequencies are shown in the lower right corner of each panel.
Trotterization
The evolution of $\hat{H}_{SO}$ in Eq. (2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization. After conjugating $\hat{H}_{SO}$ with a Hadamard gate on the qubit, the Hamiltonian can be described in the interaction picture as
$\hat{H}_{I}(t)=\frac{\epsilon}{2}\hat{\sigma}_{\Delta}(t)+\frac{\kappa}{2}\hat{\sigma}_{\Delta}(t)(\hat{b}e^{-i\nu t}+\hat{b}^{\dagger}e^{i\nu t}),$ (5)
where $\hat{\sigma}_{\Delta}(t)\equiv\hat{\sigma}_{+}e^{i\Delta t}+\hat{\sigma}_{-}e^{-i\Delta t}$ and $\hat{\sigma}_{+}$ ($\hat{\sigma}_{-}$) is the raising (lowering) operator of the qubit. The evolution with respect to $\hat{H}_{I}(t)$ is discretized into $N$ time steps, each consisting of a single-qubit rotation and a SDK operation that simulate the first and second term, respectively. Specifically, for the SDK operation at the $j$-th time step, the spin (motion) phase of the laser beams Lee et al. (2005) is set as $\Delta t_{j}$ ($\nu t_{j}$), where $t_{j}\equiv(j-\frac{1}{2})T/N\tau$ and $\tau$ is the duration of the operation for each time step. This is equivalent to setting the symmetric (anti-symmetric) detuning from the sideband transitions, referred to as the spin (motion) detuning, as $\Delta T/N\tau$ ($\nu T/N\tau$) at all time steps. This method can be straightforwardly extended to the case where multiple modes are used; however, the laser phase needs to be tuned to track the phase of each motional mode correctly sup . This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and vibrational modes Kang et al. (2024).
Randomized operations
Randomizing the control parameters is a key ingredient in our approach to simulating dissipation mechanisms in spin-boson models. By taking an average of many coherent evolutions each involving a stochastically varying parameter, we realize both heating and dephasing as described by the Lindblad master equation Cai and Barthel (2013); Chenu et al. (2017). Heating (dephasing) is used to tune the temperature (spectral linewidth) of the bath.
First, the bosonic modes at finite temperature is created by modeling heating process where the modes interact with an infinite-temperature external bath for a finite duration, described by aa pair of Lindblad operators $\hat{L}_{1}=\sqrt{\Gamma^{\prime}}\hat{b}$ and $\hat{L}_{2}=\sqrt{\Gamma^{\prime}}\hat{b}^{\dagger}$. We apply $N^{\prime}$ resonant ($\delta_{m}=0$) SDK operations of duration $\tau^{\prime}=4\Gamma^{\prime}/\tilde{\Omega}^{2}$, spin phase $\phi_{s}=0$, and motion phase $\phi_{m}$ randomly drawn from $[0,2\pi)$ for each operation. These stochastic operations are applied on the initial qubit-mode composite state |$\ket{+,n=0}$; note that $\ket{+}$ is an eigenstate of the spin operator with $\phi_{s}=0$, and thus the qubit is decoupled from these operations. When averaged over many trials of coherent evolutions, each executed with a distinct set of $N^{\prime}$ random $\phi_{m}$ values, this procedure prepares the thermal state with an average phonon number $\bar{n}=N^{\prime}\Gamma^{\prime}\tau^{\prime}$. This can be intuitively understood as the thermal state represented as an ensemble of randomly displaced coherent states (derivation provided in Supplementary Material sup ). For the experimental data in Fig. 1a and Fig. 5b, the average phonon number $\bar{n}$ of the initial thermal state is tuned by varying the duration $\tau^{\prime}$ of SDK operations, where the number of steps $N^{\prime}$ is fixed to 5. The errors in $\bar{n}$ due to finite number of steps is discussed in the Supplementary Material sup .
Next, dephasing during the time evolution, which is described by a single Lindblad operator $\hat{L}_{1}=\sqrt{\Gamma}\hat{b}$, is induced by giving random offsets to the motion detuning $\delta_{m}$ of the $N$ Trotterized SDK operations. Specifically, an offset to $\delta_{m}$ is assigned for each of the $N$ steps, with values drawn from a normal distribution characterized by a mean of zero and a standard deviation of $\sqrt{\Gamma T/N}/\tau$ sup . The qubit population dynamics averaged over many trials of coherent evolutions, each performed with a set of $N$ random $\delta_{m}$ values, give the dynamics with respect to the dephased spin-oscillator model. The entire instruction of operations is shown in Fig. 6d.
In practice, the dephasing strength $\Gamma$ is varied by tuning the standard deviation of the random offsets to $\delta_{m}$ at each Trotterization step. This enables tuning the FWHM of each Lorentzian peak in the bath’s spectral density. The errors in the simulated dephasing strength due to finite number of Trotterization steps is discussed in the Supplementary Material sup .