### v3

# Quantum Simulation of Spin-Boson Models with Structured Bath

Abstract
The spin-boson model, involving spins interacting with a bath of quantum harmonic oscillators, is a widely used representation of open quantum systems that describe many dissipative processes in physical, chemical and biological systems. Trapped ions present an ideal platform for simulating the quantum dynamics of such models, by accessing both the high-quality internal qubit states and the motional modes of the ions for spins and bosons, respectively. We demonstrate a fully programmable method to simulate dissipative dynamics of spin-boson models using a chain of trapped ions, where the initial temperature and the spectral densities of the boson bath are engineered by controlling the state of the motional modes and their coupling with qubit states. Our method provides a versatile and precise experimental tool for studying open quantum systems.
A new method to simulate open quantum systems made of spins and harmonic oscillators where the bath properties are engineered.
Interaction of a quantum system with its environment in an open dissipative setting can significantly influence its dynamics. A paradigmatic model of non-Markovian open quantum systems that captures the quantum nature of both the system and the environment is the spin-boson model Leggett et al. (1987), which is ubiquitous in the study of condensed-matter physics Caldeira and Leggett (1983); Weiss et al. (1987), chemical reactions Garg et al. (1985); Gilmore and McKenzie (2005), collective modes in atomic-photonic systems Strack and Sachdev (2011); Bhaseen et al. (2012), and biological light-harvesting complexes Huelga and Plenio (2013); Jang and Mennucci (2018). Experimentally realizing the dynamics of spin-boson models in a programmable fashion is an exciting challenge Mostame et al. (2012); Leppäkangas et al. (2018); Magazzù et al. (2018); Kim et al. (2022), which can lead to new quantum simulation approaches that reach beyond classically tractable limit.
Trapped ions provide a highly controllable system of spins and bosonic motional modes, and thus are a natural platform for simulating spin-boson models. The bosonic bath can be simulated using either a large number ($\gtrsim 50$) of motional modes Porras et al. (2008) or a handful of dissipative motional modes Lemmer et al. (2018); Schlawin et al. (2021); MacDonell et al. (2021); Kang et al. (2024). While significant experimental progress has been made in controlling the qubits and motional modes together Chen et al. (2023a); Flühmann and Home (2020); Jia et al. (2022), past works in this direction are limited to simulating a spin coherently coupled to up to two bosonic modes Whitlow et al. (2023); Valahu et al. (2023); Gorman et al. (2018); MacDonell et al. (2023). These studies could not consider the effects of dissipation often described by coupling of the system to a continuum of bath modes, which determines the system’s long-term behavior such as reaching a steady state.
Dissipative couplings come in two flavors: damping is when the quantum system exchanges energy with the bath (inelastic coupling) and eventually reaches thermal equilibrium with the bath, while dephasing is when the phase of the quantum system is randomized by interaction with the bath without energy exchange (elastic coupling). In this work, we perform quantum simulations of spin-boson models with programmable spectral densities using a 7-ion chain and up to 3 of its motional modes. The target spectral densities of the bath are decomposed into several Lorentzian peaks, which is characteristic of baths in real-world systems such as light-harvesting protein complexes Lee and Coker (2016); Kim et al. (2018). Our approach can readily simulate the dephasing model of dissipation, by adding randomness to the control parameters of the laser that drives the spin-boson interaction, such as frequency and phase. Using this technique, we study the dynamics of classic spin-boson models Leggett et al. (1987) and the vibration-assisted energy transfer (VAET) process Gorman et al. (2018). Our demonstration to engineer spectral density of the bath in analog quantum simulations is an important ingredient for achieving quantum advantage in simulating open quantum system models that are intractable using classical methods MacDonell et al. (2021); Kang et al. (2024).
We note that a recent study Wang et al. (2024) simulated a spin dissipatively coupled to many modes, albeit with limited tunability.
Dephased spin-oscillator model
The spin-boson model describes a spin coupled to a continuous bath of quantum harmonic oscillators. The Hamiltonian is given by
[…]
We decompose the spectral density $J(\omega)$ of a structured bath into a sum of multiple Lorentzian peaks, each of which can be represented by a dissipative harmonic oscillator Lemmer et al. (2018). In this work, we consider a spin coupled to several oscillators subject to constant dephasing (“dephased” spin-oscillator model). This “discrete” oscillator model is described by the Hamiltonian
$\hat{H}_{D}=\hat{H}_{S}+\sum_{l}\frac{\kappa_{l}}{2}\hat{\sigma}_{Z}\otimes(\hat{b}_{l}+\hat{b}_{l}^{\dagger})+\sum_{l}\nu_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$ (2)
and corresponding Lindblad operators $\hat{L}_{l}=\sqrt{\Gamma}_{l}\hat{b}_{l}^{\dagger}\hat{b}_{l}$, where $\kappa_{l}$ represents the coupling strength between the spin and $l$-th oscillator, $\nu_{l}$ and $\Gamma_{l}$ denote the frequency and the dephasing rate of the $l$-th oscillator, respectively, and $\hat{b}_{l}$ is the annihilation operator of the $l$-th oscillator. The composite state $\rho$ of spin and oscillators follows the Lindblad master equation $\dot{\rho}=-i[\hat{H}_{D},\rho]+\sum_{l}(\hat{L}_{l}\rho\hat{L}_{l}^{\dagger}-\frac{1}{2}\{\hat{L}_{l}^{\dagger}\hat{L}_{l},\rho\})$. In a reasonable regime of parameters ($\Gamma_{l}<\nu_{l}/2$, $\beta\ll 2\pi(\Gamma_{l}/2)^{-1}$) Lemmer et al. (2018), this bath of dephased oscillators is assigned a spectral density composed of Lorentzian peaks sup
$J_{\rm Lo}(\omega)=\sum_{l}\kappa_{l}^{2}\Big{(}\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega-\nu_{l})^{2}}-\frac{\Gamma_{l}/2}{(\Gamma_{l}/2)^{2}+(\omega+\nu_{l})^{2}}\Big{)},$ (3)
where we assume that each oscillator is initially in the thermal state with average phonon number $\bar{n}(\nu_{l})$.
The spectral density of the bath determines the dynamics of the spin to leading order in $\bar{\kappa}T$, where $\bar{\kappa}$ is the average coupling strength and $T$ is the evolution time sup . Therefore, the spin dynamics of the dephased spin-oscillator model approximately match those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ in the weak-$\bar{\kappa}$ regime. In the unbiased ($\epsilon=0$) spin-boson model Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018), the equilibrium donor-state population is $1/2$ for both models, and we expect the population dynamics of the two models show an excellent match at all times for weak spin-bath coupling. For the case of biased spin-boson models in the stronger coupling regime ($\epsilon\neq 0$, such as discussed in the VAET case), the dynamics of the dephased spin-oscillator model we simulate can deviate significantly from the traditional spin-boson model.
Figure 1: Schematic of the experimental setup. The ions are confined in a micro-fabricated surface trap. The spin states are encoded as the hyperfine clock states (qubit states) and the bosonic bath modes are encoded as the collective radial motional modes of the ion chain. We strategically use the center ion and the three symmetric modes (except for the center-of-mass mode, as shown in the upper left corner) in the experiment to minimize cross-mode couplings sup . The laser frequencies for the SDK operations are determined by the frequency difference between the two Raman beams, and are depicted by the red and blue lines on the frequency axis detuned by $\delta_{m}$ from the mode frequency $\omega_{m}$, with the color gradients indicating the random frequencies used to implement the decoherence simulation.
A more common model for dissipation is a spin coupled to oscillators subject to constant damping (“damped” spin-oscillator model) described by a pair of Lindblad operators $\hat{L}_{l1}=\sqrt{\Gamma_{l}(\bar{n}(\nu_{l})+1)}\hat{b}_{l}$ and $\hat{L}_{l2}=\sqrt{\Gamma_{l}\bar{n}(\nu_{l})}\hat{b}_{l}^{\dagger}$ Lemmer et al. (2018); Schlawin et al. (2021), where $\Gamma_{l}$ here is the damping rate of the $l$-th oscillator. A stronger result holds for this model: the spin dynamics are non-perturbatively equivalent to those of the spin-boson model with the same spectral density, up to all orders in $\bar{\kappa}T$ Tamascelli et al. (2018); sup . Damped oscillators can be realized in trapped ions using sympathetic cooling on a chain of ions with multiple atomic species or isotopes Lemmer et al. (2018); Schlawin et al. (2021). The dephased oscillator model developed in this work expands the ability to simulate a broader set of bath models not experimentally accessible before.
Experimental Implementation
Figure 2: Demonstration of the tunability of the bath’s initial temperature and spectral linewidth. (a) Dynamics of the coherent spin-single oscillator model $\hat{H}_{D}$ with various values of average phonon number $\bar{n}$ for the oscillator’s initial state. (b) Dynamics of the model with $\bar{n}=0$, $J(\omega)=J_{\rm Lo}(\omega)$ (single peak) and various values of $\Gamma_{1}\equiv\Gamma$. For both (a) and (b), $\epsilon=0$, $\nu_{1}=\Delta$, and $\kappa_{1}=0.1\Delta$. Lines and dots represent theoretical predictions and experimental data, respectively. The error bars (size comparable to the symbols in these two plots) denote measured standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times. (c) Revival amplitude versus $\bar{n}$. Predictions of the original (solid line) model and the noise-aware (dashed line) model are compared with the experimental results (circles). (d) Measured values of $\bar{n}$ fitted to the predictions of the noise-aware model. Solid line depicts $y=x+\bar{n}_{0}$, where $\bar{n}_{0}=0.036$ is the expected offset in the phonon number due to imperfect cooling. (e) Measured values of $\Gamma$ fitted to the theoretical predictions with $\bar{n}=\bar{n}_{0}$. Solid line depicts $y=x$. In (d) and (e), the blue and brown shaded regions represent the expected uncertainty of the value when the populations are averaged over 20 and 200 random trials, respectively. (f) Lorentzian spectral densities for various target (solid) and fitted (dashed) values of $\Gamma$. The shaded region represents the expected uncertainty of $\Gamma$ when averaged over 20 random trials.
The simulations of spin-boson models are performed using a linear chain of seven 171Yb+ ions (Figure 1). Two hyperfine internal states (|$\ket{0}$ and |$\ket{1}$) are used to represent the spin (or qubit). The ions sit 68 $\mu$m above the surface of the trap Revelle (2020), and the heating rate and decoherence time of the zig-zag motional mode are $3.6(3)$ quanta/s and $5.2(7)$ ms, respectively. The other motional modes used have similar magnitudes. Details of our experimental setup can be found in Ref. Wang et al. (2020). Using Doppler, electromagnetically-induced transparency (EIT), and sideband cooling techniques, the zig-zag motional mode can be cooled to near ground state with $\bar{n}=0.036(16)$. Standard qubit manipulation techniques can be used to initialize and measure the qubits, and single-qubit operations are driven by stimulated Raman transitions. The spin-oscillator coupling is simulated using the spin-dependent kick (SDK) operation induced by simultaneous application of blue- and red-sideband transitions sup .
The evolution of $\hat{H}_{D}$ in Eq. (2) is mapped to a sequence of single-qubit rotations and SDK operations via Trotterization. We apply time evolution operator over a short time interval corresponding to each term in the Hamiltonian in interaction picture, and repeat them over many time intervals to simulate the evolution dynamics sup . This approach can be readily extended to simulating more complicated molecular models consisting of many electronic states and bath modes Sun et al. (2023); Kang et al. (2024).
The dephased spin-oscillator model can be simulated using trapped ions by applying the SDK operations, which induce the coupling between the qubit and the motional mode. By adding randomness to the control parameters of the SDK operations and averaging the results over many random trials, we can implement dissipative processes such as preparing the thermal state and simulating the dephasing of the motional mode. The key idea is that certain dissipative evolutions described by the Lindblad master equation are equivalent to an average of many coherent evolutions, each subject to a stochastic Hamiltonian Chenu et al. (2017). Details of the procedure is described in Supplementary Material sup .
Single-oscillator case
We first consider a single-oscillator case ($l=1$), and demonstrate the programmability of the bath’s initial temperature quantified by the oscillator’s average phonon number $\bar{n}$. Controlled amounts of resonant SDK operations with stochastically varying phase are applied to the ion prepared near the motional ground state, which prepares the thermal state equivalent to an ensemble of randomly displaced coherent states sup . Figure 2a illustrates the simulated time evolution of the donor state population when it is coupled to such initial bath state, for various values of $\bar{n}$. The dynamics exhibit coherent oscillations with an envelope of slow collapse and revival (“beat-note”) that signifies resonant ($\nu_{1}=\Delta$) energy exchange between the system and the bath oscillator Eberly et al. (1980). A thermal state of larger $\bar{n}$ is a mixture of more “scattered” coherent states, leading to reduced beat-note amplitude at $t\approx 2\pi/\Delta\times 10.5$.
We compare our experimental results with the predictions of noise-aware models, reflecting the noise processes in the experimental setup. We fit the experimental data using an oscillator model with initial phonon number $\bar{n}+\bar{n}_{0}$ and Lorentzian spectral density with full width at half maximum (FWHM) $\Gamma_{1}\equiv\Gamma$ [Eq. (3)], where the values $\bar{n}_{0}=0.036$ and $\Gamma=0.0022\Delta$ are extracted from independent measurements characterizing imperfect cooling and finite motional coherence time, respectively. The experimentally measured revival amplitude (Fig. 2c) and fitted $\bar{n}$ (Fig. 2d) match well with the predictions of noise-aware models. This shows that well-characterized experimental noise can serve as the baseline values for the thermal excitation and dephasing rate in the open quantum system model, and thus is not always a bottleneck for the accuracy of our simulations.
Next, we show the tunability of the dephasing rate $\Gamma$, which corresponds to the linewidth of the spin-boson model’s spectral density. Dephasing is implemented by adding random frequency offsets to the SDK operations that simulate the time evolution sup . This mimics the random fluctuations of the motional-mode frequency, following the stochastic model of motional dephasing noise in trapped-ion systems Wang et al. (2020). Figure 2b illustrates the dynamics for various values of $\Gamma$ between 0 and $0.4\Delta$, which exhibit a crossover from underdamped to overdamped dynamics as $\Gamma$ is increased (critical damping condition at $\Gamma\approx 0.18\Delta$). As the damping is increased beyond the critical point, the beat-note that signifies the excitation transfer from the spin to boson back to spin is suppressed, and there is a monotonic decrease in the oscillation of the donor population in the overdamped regime. We notice a $\pi$ phase flip in the oscillation of the donor population in the beat-note in the underdamped case, signified by the phase difference between the underdamped and overdamped cases at $t\approx 2\pi/\Delta\times 10$.
Figure 3: Simulations of molecular energy transfer in model structured baths. (a) Simulated spectral density of the spin-boson model’s bath, where the electronic coupling strength $\Delta=500\>\text{cm}^{-1}$. The orange and brown solid curves illustrate the spectral densities summed over two and three peaks, respectively. (b) Dashed lines represent the radial motional modes that are coupled by the Raman beam. Black or grey dashed lines represent modes with non-zero or zero coupling strength to the target ion, respectively. The three solid lines and the shaded regions depict the mean value and the standard deviation of the randomized pulse frequencies, and the color corresponds to the Lorentzian peak generating the final spectral density in (a). (c) Time evolution of the spin-boson models with $J(\omega)$ in (a). The bath temperature is 77 K such that $\bar{n}\approx 0.1$, which is set for trapped ions’ motional modes using imperfect sideband cooling. Curves and dots represent theoretical predictions and experimental data, respectively. Error bars are derived as in Fig. 2.
The uncertainty of $\bar{n}$ and $\Gamma$ realized in the experiment can be reduced by performing a larger number of random trials, each trial using a different set of random parameters. The shaded areas in Fig. 2d and e show the standard deviation of fitted values for $\bar{n}$ and $\Gamma$, obtained by numerical simulation of results averaged over 20 (pink) or 200 (blue) random trials. Our experiments are limited to 20 random trials due to the compiling delay required for each set of control parameters, and each trial is repeated 100 times to obtain the expectation value for the donor population. The experimental data is consistent with the standard deviation derived from 20 numerical trials. We expect to readily increase the number of random trials with fewer repetitions each in future experiments with faster compilation tools Dalvi et al. (2023); sup .
Engineering bath spectral densities
Figure 4: Simulations based on Leggett’s spin-boson model. (a) Spectral densities of the Leggett’s model. Dot-dashed curves are the sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities described by Eq. (4) ($A=0.1$, $\omega_{c}=10\Delta$). Solid curves are the spectral densities simulated in our experiments using two motional modes. These spectral densities are designed to match the dot-dashed curves near $\omega=\Delta$. Insets compare the solid and dot-dashed curves within the interval $\omega\in[0.85\Delta,1.15\Delta]$. (b) (c) (d) Dynamics of the spin-boson model for $s=0.5$, $1.0$, and $2.0$, respectively. The solid curves represent the theoretical predictions of the spin-boson models with spectral densities given by the solid curves in (a). The red dashed curves are expected results, derived from numerical simulations of the density matrix of the qubit and motional modes subject to control operations and hardware noise. Circles denote experimental data and error bars indicate the standard deviation over 60, 20, and 20 trials for (b), (c), and (d), respectively.
We can use multiple oscillator modes to engineer a target spectral density structure for the oscillator bath. We simulate a bath of spectral density composed of up to 3 Lorentzian peaks, each represented by a motional mode of the ion chain. The bath spectral density $J(\omega)$ in Fig. 3a is simulated using the SDK operations over two or three motional modes. Figure 3b shows the frequencies of the laser pulses used to drive the SDK, defined as the detuning of laser beams from the carrier transition frequency. The 1st, 3rd, and 5th motional modes are used, and the detuning of the laser frequency from each motional mode determines the frequency at which the bath spectral density is considered in Fig. 3a. We set the electronic coupling strength $\Delta=500$ cm-1 Chenu and Scholes (2015), such that at temperature 77 K, $\bar{n}\approx 0.1$ for the near-resonance vibrational modes.
Figure 3c shows the results for simulating the dynamics of spin-boson models with spectral densities composed of 2 and 3 Lorentzian peaks, respectively. The theoretical predictions are obtained using the time-dependent density-matrix renormalization group algorithm in the interaction picture Nuomin et al. (2022). The donor population features underdamped coherent oscillations, with contributions from all three modes. The measured population closely follows the theoretical predictions.
The coherence time of the motional modes in our setup of $\sim 5.2(7)$ ms is the leading source of noise Wang et al. (2020); Kang et al. (2023a), and is comparable to the experimental time for simulating the time evolution under spectral densities generated from 2 mode (1.63 ms) and 3-mode (2.56 ms) cases. For these simulations, the contribution of motional decoherence in the experiment is identical to the random kicks that we apply in the spin-oscillator model to simulate the bath, and can be incorporated as the source of dephasing in our simulation with proper calibration Kang et al. (2023a).
Leggett spin-boson models
We further simulate the spin-boson models inspired by Leggett et al. Leggett et al. (1987). The spectral density of the bath is given by
$J_{\rm Legg}(\omega)=A\omega^{s}\omega_{c}^{1-s}e^{-\omega/\omega_{c}},$ (4)
where $\omega_{c}$ is the cutoff frequency, and $s<1$, $s=1$, and $s>1$ represent the sub-Ohmic, Ohmic and super-Ohmic baths, respectively. This model is widely used to characterize the noise in nanomechanical devices Galland et al. (2008), superconducting circuits Magazzù et al. (2018); Kaur et al. (2021), and proton-transfer reactions Ohta et al. (2006); Craig et al. (2007). The spin dynamics are expected to exhibit a phase transition from coherent oscillations to incoherent decay or localization as $s$ is decreased or the amplitude $A$ is increased, which has been a subject of longstanding research using analytical tools Leggett et al. (1987); Bulla et al. (2003) and classical simulations Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Nalbach and Thorwart (2013).
Recognizing that the spin exchanges energy with the bath only near its resonance, we use a sum of Lorentzian lines from motional modes to approximate the spectral density $J_{\rm Legg}(\omega)$ near the resonance frequency $\Delta$. We utilize well-established global optimization algorithms to find the optimal set of parameters ($\kappa_{l}$, $\Gamma_{l}$, and $\nu_{l}$ in Eq. (3)) that best represent $J_{\rm Legg}(\omega)$ sup . Using this approach, we consider weak spin-bath coupling ($A=0.1$) and match the spectral density within the target bandwidth $\omega\in[0.9\Delta,1.1\Delta]$ with 2 Lorentzian peaks. Figure 4a shows the obtained spectra from the modes (solid lines) that approximate sub-Ohmic ($s=0.5$), Ohmic ($s=1.0$), and super-Ohmic ($s=2.0$) spectral densities (dot-dashed lines), respectively. The results for the donor population dynamics interacting with a bath featuring these approximated spectral densities are shown in Fig. 4b-d. We observe a crossover from near-coherent ($s=2$) to fully damped ($s=0.5$) oscillations in the simulated time scale, owing to the different levels of spectral density $J(\omega=\Delta)$ near resonance. The observed behavior qualitatively agrees with the predictions of the Leggett model: for super-Ohmic case, coherence is maintained due to low spectral density near resonance, while for sub-Ohmic case, large spectral density induces strong damping to the point where the system is localized. We note that the deviations in the $s=2$ case in Fig. 4d result from insufficient number of Trotterization steps sup .
Vibration-assisted energy transfer
Finally, we simulate the VAET model Gorman et al. (2018) with nonzero energy detuning $\epsilon$ between the spin states. Here we set the spin parameters as $\Delta=30\>\text{cm}^{-1}$ and $\epsilon=100\>\text{cm}^{-1}$. In the absence of coupling to a vibrational mode ($\kappa=0\>\text{cm}^{-1}$), energy transfer between the detuned spin states is suppressed; however, for nonzero $\kappa$, the energy transfer can be activated if the mode frequency $\nu$ meets the resonance condition (single mode’s index $l=1$ omitted).
Figure 5a shows the experimental data for simulating the coherent spin-single oscillator model $\hat{H}_{D}$ with a relatively large coupling strength ($\kappa=30\>\text{cm}^{-1}$). When $\nu$ satisfies the resonance condition of VAET ($\nu=\sqrt{\Delta^{2}+\epsilon^{2}}/k\approx 104\>\text{cm}^{-1}/k$ where $k$ is an integer), energy transfer occurs Gorman et al. (2018). Otherwise, energy remains localized at the donor state. Figure 5b shows the impact of initial temperature of the oscillator on the resonant VAET ($\nu=104\>\text{cm}^{-1}$). We create baths of two temperatures 0 K and 300 K corresponding to initial average phonon number $\bar{n}=0$ and $2.0$, respectively sup . The energy transfer is faster for higher $\bar{n}$ in the beginning as more vibrational quanta assist the energy transfer, but it also leads to a rapid decay of oscillations over time as the coherence is lost.
In Fig. 5c, we simulate the dephased spin-oscillator model with $\Gamma=10\>\text{cm}^{-1}$ and various values of $\kappa$, to study the impact of dissipative coupling. We note that in the weak-$\kappa$ regime (or short evolution time), the results agree with the theoretical predictions of the damped spin-oscillator model, which is equivalent to those of the spin-boson model with $J(\omega)=J_{\rm Lo}(\omega)$ Lemmer et al. (2018); Tamascelli et al. (2018); however, as $\kappa$ is increased, the dephased and damped models’ dynamics deviate. This is expected because unlike the damped model, the dephased model does not extract energy from the spin-oscillator system, so the donor population will not decay towards zero even after a long time evolution when coupled to a single-oscillator bath. This example demonstrates that our method provides the ability to simulate a wider range of dissipative channels in the study of open system dynamics.
Figure 5: Simulations of vibration-assisted energy transfer (VAET). We set $\epsilon=100\>\text{cm}^{-1}$ and $\Delta=30\>\text{cm}^{-1}$. (a) (b) Time evolution of the donor-state population for coherent VAET models with $\kappa=30\>\text{cm}^{-1}$. Solid and dashed curves depict theoretical predictions of the ideal target and the noise-aware model, respectively. Circles represent experimental data and error bars are standard deviation over 100 repetitions. In (a), bath temperature is fixed at 0 K ($\bar{n}=0$) and $\nu$ is varied. In (b), $\nu$ is fixed to the resonant value $104\>\text{cm}^{-1}$ and two temperatures 0 K and 300 K, which correspond to $\bar{n}=0$ and 2.0, respectively, are simulated. (c) Time evolution of the donor-state population for dissipative VAET models with $\nu=104\>\text{cm}^{-1}$ and $\Gamma=10\>\text{cm}^{-1}$ at 0 K. Solid and dot-dashed curves depict theoretical predictions of the dephased spin-oscillator model and the damped spin-oscillator model, respectively. Circles denote experimental data. For $\kappa=0\>\text{cm}^{-1}$, error bars indicate the standard deviation over 100 repetitions. For $\kappa=10$ and $20\>\text{cm}^{-1}$, error bars indicate the standard deviation over 20 trials, each trial performed with randomly drawn set of parameters and repeated 100 times.
Discussions and outlook
While classical simulation approaches to open dissipative systems continue to improve with advancements in numerical methods, they encounter fundamental challenges when dealing with large quantum systems that span exponentially large Hilbert space MacDonell et al. (2021); Kang et al. (2024). Take, for example, the Dissipation-Assisted Matrix Product Factorization (DAMPF) method, a state-of-the-art method for simulating open quantum systems Somoza et al. (2019). For simulating a one-dimensional chain of $N_{s}$ electronic states each coupled to a bath of $N_{b}$ damped bosonic modes, the computational complexity of DAMPF scales as $\mathcal{O}(N_{s}^{5}N_{b}D_{b}^{2}\chi^{3})$, where $D_{b}$ is the number of energy levels that represent each bosonic mode and $\chi$ is the bond dimension in the matrix product operator representation of the system-bath density matrix. A critical issue is the rapid, potentially exponential increase in $\chi$ with respect to the system-bath coupling strength, for a fixed target accuracy in the population dynamics. For these types of problems, trapped-ion simulators may offer a more scalable alternative, as the number of operations scales only as $\mathcal{O}(N_{s}N_{b})$ and the total duration of SDK operations grows linearly with the system-bath coupling strength Kang et al. (2024). The current state-of-the-art trapped-ion systems offer several dozens of highly coherent spins and bosonic modes that are very densely coupled with each other, the complete quantum description of which is totally intractable with classical computers of today Cetina et al. (2022); Chen et al. (2023b); DeCross et al. (2024). Developing various control methods over these quantum systems to study important physical, chemical, biological and material systems of critical interest will open up a new computational framework in science.
The methods developed in this work utilize fully programmable control methods to simulate dissipative processes. The dephased oscillator method can be used to realize non-Gaussian bath models Barthel and Zhang (2022), relevant for studying physical processes such as phase transitions in strongly-coupled spin-boson models Leggett et al. (1987); Wang and Thoss (2008); Duan et al. (2017); Strathearn et al. (2018); Nalbach and Thorwart (2013) and role of bosonic dephasing in molecular energy transfer Stock and Domcke (1990); Banin et al. (1994). The capability to measure the probability distribution of the motional modes Flühmann and Home (2020); Jia et al. (2022); Valahu et al. (2023); Whitlow et al. (2023) may further reveal the effects of intricate vibronic interactions in biological and chemical reaction dynamics. High accuracy simulations will require development of efficient characterization methods Kang et al. (2023b) and novel control protocols in the trapped-ion system sup ; Lobser et al. (2023) to implement various interaction terms in the target Hamiltonian. Combined with damped spin-boson models that can be implemented using sympathetic cooling Lemmer et al. (2018); Schlawin et al. (2021), the long-chain trapped-ion systems can provide a versatile platform for simulating dynamics of complex open quantum systems.
During the preparation of our manuscript we became aware of a complementary work So et al. (2024) which simulates the electron-transfer model using sympathetic cooling on the trapped ions’ motional mode.
Acknowledgments: The authors thank Yichao Yu for his valuable discussions about the phase tracking method, Daniel Lobser for his valuable suggestion in controlling the RFSoC control software, and Thomas Barthel for his course on open quantum systems that inspired this work. Funding: This research is funded by the Office of the Director of National Intelligence - Intelligence Advanced Research Projects Activity, through the ARO contract W911NF-16-1-0082 (provided the experimental apparatus used in the demonstration), the DOE BES award DE-SC0019400 (theoretical modeling), and the NSF Quantum Leap Challenge Institute for Robust Quantum Simulation Grant No. OMA-2120757 (experimental concepts and validation). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (data analysis and validation). Author contributions: K.S., M.K., K.R.B. and J.K. conceived the idea and developed detailed methodology for experiments and modeling, K.S. and G.S. performed the experiments, M.K., H.N. and D.N.B performed modeling and simulations, K.S., M.K., K.R.B. and J.K. analyzed and validated the data, K.R.B. and J.K. supervised the project, K.S. and M.K. wrote the original draft, and K.S., M.K. and J.K. revised the manuscript. All authors discussed the results and reviewed the final manuscript. Competing Interests: J.K. and K.R.B are scientific advisors to IonQ, Inc., and J.K. is a shareholder of IonQ, Inc. Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or Supplementary Materials. We will provide raw data to anyone who might be interested.